# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v2

@@ V-B2 Backward Message Passing
VI Simulation Results
In this section, we evaluate the performance of the proposed channel estimation scheme under various system setups. The simulation parameters are shown in Table II. In particular, we consider normalized mean square error (NMSE) as performance metrics, which is defined as $\mathrm{NMSE}\triangleq{\lVert\hat{\mathbf{H}}-\mathbf{H}\rVert^{2}_{\mathrm{F}}}/{\lVert\mathbf{H}\lVert^{2}_{\mathrm{F}}}$, where $\mathbf{H}$ and $\hat{\mathbf{H}}$ are the true channel and estimated channel, respectively. In addition, the SNR is defined in received side, which is given by $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$. Additionally, we compare the proposed MMV-VMP algorithm with the following baselines:
@@ VI-A Convergence and Complexity Performance
Figure 4: Convergence behavior of MMV-VMP.
Fig. 4 illustrates the convergence behavior of the proposed MMV-VMP algorithm by plotting the NMSE against the number of iterations under various pilot lengths and SNR conditions. The results clearly demonstrate a consistent monotonic decrease in NMSE across all iterations, indicating the algorithm’s stable and reliable convergence across different simulation scenarios. Based on the trade-off between estimation accuracy and computational complexity, it is observed that the NMSE stabilizes after approximately 20 iterations. Therefore, for subsequent simulations, the maximum number of iterations can be effectively limited to 20 without compromising performance.
Figure 5: Complexity evaluation of different algorithms.
Fig. 5 presents a computational complexity analysis by comparing the CPU runtime of various algorithms. Among these, SOMP exhibits the shortest runtime. This efficiency is attributed to its computational simplicity, involving only matrix-vector product between the residual and the measurement matrix, as well as a projection operation between the received signal and the low-dimensional basis matrix. In contrast, SBL-based methods such as StdSBL, PC-SBL, and VSP require significantly more computational time. This is primarily due to the matrix inversion operations inherent in their implementation, which increase computational complexity. In contrast, the proposed MMV-VMP algorithm demonstrates a shorter runtime compared to StdSBL, PC-SBL, and VSP due to avoidance of matrix inversion operations and stable convergence properties.
@@ VI-B NMSE versus SNR and Pilot Symbol Number
Fig. 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$. The results reveal that algorithms such as SOMP [29], StdSBL [36], and UAMP-SBL [33], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-VMP. Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity. However, it does not consider global block sparsity. In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity. The proposed MMV-VMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both global block sparsity and local common-delay sparsity. This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
Figure 7: NMSE versus pilot symbol number.
Fig. 7 evaluates the estimation performance of various algorithms under different compression ratios by varying the number of pilot symbols, $P$, with the fixed SNR of 2 dB. As the number of pilot symbols varies from 6 to 12, the compression ratio $M/N$ ranges from 0.375 to 0.75. We can obtain two key observations: 1) Across most compression ratios, the proposed MMV-VMP algorithm outperforms baseline methods. This result highlights its robust sparse recovery capability under varying compression conditions. Moreover, the superior performance demonstrates that the proposed algorithm is a low-overhead solution, requiring fewer pilot symbols to achieve comparable or better performance than other algorithms. 2) As the pilot length increases, the rate of performance improvement gradually diminishes, eventually approaching saturation. This observation suggests that selecting a moderate value of $P$ is sufficient to strike a balance between estimation performance and pilot overhead.
@@ VI-C NMSE versus Path Number
Fig. 8 illustrates the NMSE performance of various algorithms as a function of the number of propagation paths, with $P=8$ and $\mathrm{SNR}=2$dB. The SOMP algorithm exhibits significant sensitivity to prior knowledge of the number of paths, with its performance degrading sharply when the assumed and actual path numbers differ. In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios. Notably, the proposed MMV-VMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
VI-D Ablation Study to Evaluate the Effects of Prior Models
@@
To evaluate the effectiveness of the proposed column-wise prior model, the MMV-VMP algorithm is compared with two baselines: (1) a column-wise hierarchical prior without the MRF, labeled as “Column-Wise w.o. MRF,” and (2) a column-wise prior without the common-delay sparsity mechanism, labeled as “Column-Wise w.o. CD.” Fig. 9 reveals two key observations: 1) Both the MRF and the common-delay sparsity mechanism contribute to improved estimation performance. Notably, the improvement introduced by the MRF is more significant than that of the common-delay sparsity mechanism. 2) The simultaneous use of the MRF and the common-delay sparsity mechanism better captures the structured sparsity of the channel, leading to further enhancement in estimation accuracy.
To showcase the estimation performance of different prior more intuitively, Fig. 10 provides the results of a Monte Carlo simulation with $\mathrm{SNR}=2$ dB and $P=8$. The original spatial-delay channels are compared with the reconstructed channels obtained using “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and the proposed MMV-VMP algorithm. The figure clearly shows that the proposed MMV-VMP algorithm achieves the best visual quality, accurately reconstructing the spatial-delay channels and demonstrating its superior capability in sparse recovery. As expected, “Column-Wise w.o. MRF” provides the poorest reconstruction quality due to its limited ability to capture global block sparsity. Moreover, compared to “Column-Wise w.o. CD,” the reconstructed channels by MMV-VMP exhibit reduced power dispersion phenomenon near the non-zero regions, attributed to the incorporation of the common-delay sparsity mechanism.
@@
(d) MMV-VMP.
Figure 10: Original and reconstructed channels by “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and MMV-VMP algorithm.
@@ VII Conclusions
In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, spatially non-stationary (SnS) properties, and dual-wideband effects. We first rigorously quantified the angular and delay spread effects of the SnS dual-wideband XL-MIMO channel in the angular-delay domain. Building upon this, we highlighted the global and local common-delay sparsity inherent in the angular-delay channels. Leveraging this structured sparsity, we proposed a computationally efficient MMV-VMP algorithm, where variational message passing is employed to handle the complex factor graph arising from the column-wise hierarchical prior. Simulation results demonstrate that the proposed MMV-VMP algorithm consistently outperforms existing methods, offering robust performance with moderate computational complexity. Additionally, an ablation study was conducted to assess the effectiveness of the column-wise hierarchical prior, further validating the advantages of the proposed approach.
@@ Appendix A Proof of Lemma 1
The proof of Theorem 1 is portioned as two part. We first prove that the angular-delay matrix of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ is block-sparse. Then, applying the shift property of 2D-IDFT, the block sparsity of $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}$ can be obtained. For notation simplification, we discard the subscript $l$, $\varphi_{l}$, and $\psi_{l}$ in the subsequent derivations. We begin by examining the angular-domain transformation of $\boldsymbol{\Theta}$ with $\cos\vartheta>0$. Mathematically, $\boldsymbol{\Theta}$ can be rewritten in a chirp signal form as
$[\boldsymbol{\Theta}]_{n,k}=\mathrm{e}^{-\mathrm{j}2\pi\left(a_{k}n-\frac{b_{k}}{2}n^{2}\right)},$ (A.1)
where $a_{k}={f_{k}\psi}/{f_{c}}$ and $b_{k}={2f_{k}\varphi}/{f_{c}}$ denote the starting spatial frequency and chirp rate, respectively. Utilizing the inherent chirp properties of Eq. (A.1), the IDFT of $\boldsymbol{\Theta}$ for the fixed $k$, i.e., $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}=\mathbf{F}^{\mathrm{H}}_{\mathrm{A}}\boldsymbol{\theta}(f_{k})\in\mathbb{C}^{N\times 1}$ exhibits a distinct rectangular shape [38], where $\mathbf{F}_{\mathrm{A}}$ denotes a $N$-dimension DFT matrix. The boundary spatial frequencies are given by
$i_{s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (A.2)
$i_{e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d},$ (A.3)
where $n_{s}$ and $n_{e}$ denote the start and ending index of VR $\boldsymbol{\phi}$, i.e., $\boldsymbol{\phi}\in[n_{s},n_{e}]$. In this case, due to $a_{k}>0$ and $b_{k}>0$, we have $i_{s}(k)>i_{e}(k)$. Consequently, $[\boldsymbol{\Xi}_{\mathrm{A}}]_{:,k}$ demonstrates block sparsity, with its prominent spatial frequency components concentrated from $i_{e}(k)$ to $i_{s}(k)$.
@@
Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{e}^{\mathrm{min}}=\min_{k}i_{e}(k)$ to $i_{s}^{\mathrm{max}}=\max_{k}i_{s}(k)$. As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
$\boldsymbol{\Xi}_{\mathrm{A}}=\left[\mathbf{0}_{K\times(I_{e}-1)}^{\mathrm{T}},\mathbf{U}_{K\times(I_{s}-I_{e}+1)}^{\mathrm{T}},\mathbf{0}_{K\times(N-I_{s})}^{\mathrm{T}}\right]^{\mathrm{T}},$ (A.4)
@@
where $\mathbf{U}$ is a non-zero matrix; index $I_{e}$ and $I_{s}$ are given by $I_{e}=\lceil{(i_{e}^{\mathrm{min}}+1)N}/{2}\rceil$ and $I_{s}=\lceil{(i_{s}^{\mathrm{max}}+1)N}/{2}$.
@@
Further, we examine the IDFT of $\boldsymbol{\Theta}$ in the delay domain, i.e., $\boldsymbol{\Xi}_{\mathrm{D}}=\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{*}$, where $\mathbf{F}_{\mathrm{D}}$ denote a $K$-dimension DFT matrix. Specifically, we have
n,m $\propto\frac{1}{\sqrt{K}}\sum_{k=0}^{K-1}\mathrm{e}^{\mathrm{j}\frac{2\pi}{K}mk}\mathrm{e}^{-\mathrm{j}2\pi kt_{n}}$ (A.5)
$\propto\frac{1}{\sqrt{K}}\frac{\sin\left(\pi K(t_{n}-\frac{m}{K})\right)}{\sin\left(\pi(t_{n}-\frac{m}{K})\right)},$
@@
where $t_{n}=n\frac{f_{s}\psi}{Kf_{c}}-n^{2}\frac{f_{s}\varphi}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
@@
$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{n}-\frac{m}{K}\right).$ (A.6)
@@
From (A.6), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{n}K$. However, due to the finite sampling size $K$ in practice, we consider the range from $j_{e}(n)=(t_{n}-{2}/{K})K$ to $j_{s}(n)=(t_{n}+{2}/{K})K$ to collect the delay indices near the peak effectively. Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
$\boldsymbol{\Xi}_{\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{e}-1)},\mathbf{L}_{N\times(J_{s}-J_{e}+1)},\mathbf{0}_{N\times(K-J_{s})}\right],$ (A.7)
where $J_{e}=\min_{n}j_{e}(n)$ to $J_{s}=\max_{n}j_{s}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
@@
According to row-wise and column-wise block sparsity of (A.4) and (A.7), we assert that $\boldsymbol{\Xi}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{1}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{e}\leq n\leq I_{s},J_{e}\leq k\leq J_{s}\right\}$, i.e.,
@@
Next, we examine the 2D-IDFT of channel $\mathbf{H}$, $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$. By leveraging the linearity of the 2D-IDFT, we can focus solely on analyzing the $l$-th path. Define $\mathbf{H}_{l}=\alpha_{l}\left(\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})\right)\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}$, whose $(n,k)$-th element is given by
n,k $\propto\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}+f_{k}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$ (A.9)
$\overset{(a)}{\propto}\sum_{p=1}^{P}\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k},$
where $(a)$ is obtained by decomposing the chirp signal $\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}$ into the superposition of $P$ complex exponential signals with frequencies $\epsilon_{l,p}=\psi_{l}-2n_{l,e}\varphi_{l}+\eta_{l}(p-1)$ with $\eta_{l}={2N\varphi_{l}}/{P}$ and $p\in\{p\in\mathbb{N}^{+}\mid\psi_{l}-2n_{l,e}\varphi_{l}\leq\eta_{l}p\leq\psi_{l}-2n_{l,s}\varphi_{l}\}$. Furthermore, we focus on $\mathbf{H}_{l,p}(n,k)=\mathrm{e}^{-\mathrm{j}2\pi\left(\frac{n}{N}N\epsilon_{l,p}+\frac{k}{K}f_{s}\tau_{l}\right)}[\boldsymbol{\Theta}]_{n,k}$. Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{A}_{l,2,p}$ with
$\mathcal{A}_{l,2,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (A.10)
$k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l,1}\right\}.$
Here, $\mathcal{A}_{l,1}$ is the non-zero region for $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$, and mod$(a,m)$ is the modulus of $a$ for $m$. Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{A}_{l,2}\triangleq\cup_{p}\mathcal{A}_{l,2,p}$. Thus, the proof is completed.

### v3

@@ VI Simulation Results
In this section, we evaluate the performance of the proposed channel estimation scheme under various system setups. The simulation parameters are shown in Table II. In particular, we consider normalized mean square error (NMSE) as performance metrics, which is defined as $\mathrm{NMSE}\triangleq{\lVert\hat{\mathbf{H}}-\mathbf{H}\rVert^{2}_{\mathrm{F}}}/{\lVert\mathbf{H}\lVert^{2}_{\mathrm{F}}}$, where $\mathbf{H}$ and $\hat{\mathbf{H}}$ are the true channel and estimated channel, respectively. In addition, the SNR is defined in received side, which is given by $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$. Additionally, we compare the proposed MMV-HMP algorithm with the following baselines:
@@
Figure 4: Convergence behavior of MMV-HMP.
Figure 5: Complexity evaluation of different algorithms.
@@ VI-A Convergence and Complexity Performance
Fig. 4 illustrates the convergence behavior of the proposed MMV-HMP algorithm by plotting the NMSE against the number of iterations under various pilot lengths and SNR conditions. The results clearly demonstrate a consistent monotonic decrease in NMSE across all iterations, indicating the algorithm’s stable and reliable convergence across different simulation scenarios. Based on the trade-off between estimation accuracy and computational complexity, it is observed that the NMSE stabilizes after approximately 20 iterations. Therefore, for subsequent simulations, the maximum number of iterations can be effectively limited to 20 without compromising performance.
Fig. 5 presents a computational complexity analysis by comparing the central processing unit (CPU) runtime of various algorithms. Among these, SOMP exhibits the shortest runtime. This efficiency is attributed to its computational simplicity, involving only matrix-vector product between the residual and the measurement matrix, as well as a projection operation between the received signal and the low-dimensional basis matrix. In contrast, SBL-based methods such as StdSBL, PC-SBL, and VSP require significantly more computational time. This is primarily due to the matrix inversion operations inherent in their implementation, which increase computational complexity. As anticipated, the proposed MMV-HMP algorithm demonstrates a shorter runtime compared to StdSBL, PC-SBL, and VSP due to avoidance of matrix inversion operations and stable convergence properties.
@@ VI-B NMSE versus SNR and Pilot Symbol Number
Fig. 6 illustrates the NMSE performance of various algorithms as a function of SNR for $P=8$. The results reveal that algorithms such as SOMP [36], StdSBL [46], and UAMP-SBL [40], which fail to account for both global block sparsity and local common-delay sparsity, exhibit significantly poorer NMSE performance compared to algorithms like PC-SBL, VSP, and the proposed MMV-HMP. Among block-sparsity prior-based methods, the PC-SBL algorithm leverages a pattern-coupled Gaussian prior model to effectively capture local common-delay sparsity. However, it does not consider global block sparsity. In contrast, the VSP algorithm employs a MRF-based prior to model global block sparsity but lacks the ability to incorporate local common-delay sparsity. In contrast, the proposed MMV-HMP algorithm overcomes these limitations by utilizing a tailored column-wise hierarchical prior that simultaneously incorporates both MRF structure and precision parameter sharing mechanism. This comprehensive modeling approach ensures consistently superior NMSE performance across the entire SNR range under consideration.
Figure 6: NMSE versus SNR.
@@
Fig. 7 evaluates the estimation performance of various algorithms under different compression ratios by varying the number of pilot symbols, $P$, with the fixed SNR of 2 dB. As the number of pilot symbols varies from 6 to 12, the compression ratio $M/N$ ranges from 0.375 to 0.75. We can obtain two key observations: 1) Across most compression ratios, the proposed MMV-HMP algorithm outperforms baseline methods. This result highlights its robust sparse recovery capability under varying compression conditions. Moreover, the superior performance demonstrates that the proposed algorithm is a low-overhead solution, requiring fewer pilot symbols to achieve comparable or better performance than other algorithms. 2) As the pilot length increases, the rate of performance improvement gradually diminishes, eventually approaching saturation. This observation suggests that selecting a moderate value of $P$ is sufficient to strike a balance between estimation performance and pilot overhead.
@@ VI-C NMSE versus Path Number
Fig. 8 illustrates the NMSE performance of various algorithms as a function of the number of propagation paths, with $P=8$ and $\mathrm{SNR}=2$dB. The SOMP algorithm exhibits significant sensitivity to prior knowledge of the number of paths, with its performance degrading sharply when the assumed and actual path numbers differ. In contrast, Bayesian inference-based methods, leveraging sparsity-promoting prior models, adapt effectively to variations in the number of paths, maintaining consistent and robust performance across all scenarios. Notably, the proposed MMV-HMP algorithm consistently outperforms other Bayesian methods, demonstrating superior capability in handling dynamic sparse scattering environments.
@@
(d) MMV-HMP.
Figure 10: Original and reconstructed channels by “Column-Wise w.o. MRF,” “Column-Wise w.o. CD,” and MMV-HMP algorithm.
VI-D Ablation Study to Evaluate the Effects of Prior Models
To evaluate the effectiveness of the proposed column-wise hierarchical sparse prior model, we conduct an ablation study. Specifically, the proposed MMV-HMP algorithm is compared against the following two baselines:
  • Column-Wise w.o. MRF: This baseline removes the MRF structure from the proposed column-wise hierarchical sparse prior model, thereby ignoring coefficient correlations.
  • Column-Wise w.o. CD: This baseline eliminates the precision parameter sharing strategy in the proposed prior model, thus modeling each coefficient’s precision independently without promoting local clustering.
Fig. 9 investigates the NMSE performance versus SNR for different prior models. As expected, the estimation performance of the two baselines exhibits significant degradation, as they fail to fully capture both the global and local structured sparsity. This validates the necessity of jointly employing the MRF structure and the precision parameter sharing mechanism. Moreover, it is observed that the performance degradation of the “Column-Wise w.o. MRF” model is substantially greater than that of the “Column-Wise w.o. CD” model. This indicates that the improvement introduced by incorporating the MRF is more critical than the gain achieved by the precision parameter sharing mechanism.
To intuitively demonstrate the estimation performance of different priors, Fig. 10 presents the results of a Monte Carlo simulation with $\mathrm{SNR}=2$ dB and $P=8$. The original spatial-delay channels are compared with the reconstructed channels obtained using the baseline methods and the proposed MMV-HMP algorithm. The figure clearly shows that the proposed MMV-HMP algorithm achieves the best reconstruction quality, accurately recovering the spatial-delay channels and demonstrating superior sparse recovery capability. As expected, “Column-Wise w.o. MRF” yields the poorest reconstruction, with significant residual noise, mainly due to its limited ability to capture the global block sparsity. Furthermore, compared to “Column-Wise w.o. CD,” the channels reconstructed by MMV-HMP exhibit noticeably reduced power dispersion near the non-zero regions. This improvement is attributed to the incorporation of the precision parameter sharing mechanism.
@@ VII Conclusions
In this paper, we have addressed the channel estimation problem in XL-MIMO systems, taking into account the spherical wavefront effects, SnS properties, and dual-wideband effects. We began by rigorously quantifying the angular and delay spread properties of SnS dual-wideband channels in the angular-delay domain and revealed their inherent global block sparsity and local common-delay sparsity. To exploit this structured sparsity, we proposed a computationally efficient MMV-HMP algorithm. The algorithm performs message updates along the factor graph by selectively applying either the SPA or VMP rules, depending on the structure and tractability of each factor node. Simulation results demonstrate that the proposed MMV-HMP algorithm consistently outperforms existing baseline methods, achieving robust estimation performance with moderate computational complexity. Furthermore, an ablation study confirms the effectiveness of the proposed column-wise hierarchical prior, validating its contribution to improved estimation accuracy.
@@ Appendix A Proof of Lemma 1
The proof of Lemma 1 is divided into two parts. First, we demonstrate that the angular-frequency representation $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ exhibits row-wise sparsity. Next, we show that the spatial-delay representation $\boldsymbol{\Theta}(\varphi_{,}\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ exhibits column-wise sparsity. Without loss of generality, we assume that $\cos\vartheta>0$.
We begin by examining the angular transformation of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$. For a fixed $k$, the angular transformation of $\boldsymbol{\theta}(f_{k},\varphi_{l},\psi_{l})$ is given by
$[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{m,k}=\sum_{n=1}^{N}\mathrm{e}^{\mathrm{j}\frac{2\pi}{N}mn}\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n}.$ (A.1)
@@
According to (21) and (22), it can be obtained that $[\boldsymbol{\Xi}_{l,\mathrm{A}}]_{:,k}$ demonstrates block sparsity with its prominent spatial frequency components concentrated from $i_{l,e}(k)$ to $i_{l,s}(k)$. Extending this observation across all $K$ subcarriers, the overall spatial frequency range of significant components spans from $i_{l,e}^{\mathrm{min}}=\min_{k}i_{l,e}(k)$ to $i_{l,s}^{\mathrm{max}}=\max_{k}i_{l,s}(k)$. As a result, $\boldsymbol{\Xi}_{\mathrm{A}}$ is characterized as a row-wise block matrix
$\boldsymbol{\Xi}_{l,\mathrm{A}}=\left[\mathbf{0}_{K\times(I_{l,e}-1)}^{\mathrm{T}},\mathbf{U}_{K\times(I_{l,s}-I_{l,e}+1)}^{\mathrm{T}},\mathbf{0}_{K\times(N-I_{l,s})}^{\mathrm{T}}\right]^{\mathrm{T}},$ (A.2)
@@
where $\mathbf{U}$ is a non-zero matrix; index $I_{l,e}$ and $I_{l,s}$ are given by $I_{l,e}=\lceil{(i_{l,e}^{\mathrm{min}}+1)N}/{2}\rceil$ and $I_{l,s}=\lceil{(i_{l,s}^{\mathrm{max}}+1)N}/{2}\rceil$.
@@
Further, we examine the IDFT of $\boldsymbol{\Theta}(\varphi_{l},\psi_{l})$ in the delay domain, i.e., $\boldsymbol{\Xi}_{l,\mathrm{D}}=\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, where $\mathbf{F}_{\mathrm{D}}$ denote a $K$-dimension DFT matrix. Specifically, we have
n,m $\propto\frac{1}{\sqrt{K}}\sum_{k=0}^{K-1}\mathrm{e}^{\mathrm{j}\frac{2\pi}{K}mk}\mathrm{e}^{-\mathrm{j}2\pi kt_{l,n}}$ (A.3)
$\propto\frac{1}{\sqrt{K}}\frac{\sin\left(\pi K(t_{l,n}-\frac{m}{K})\right)}{\sin\left(\pi(t_{l,n}-\frac{m}{K})\right)},$
@@
where $t_{l,n}=n^{2}\frac{f_{s}\varphi_{l}}{Kf_{c}}-n\frac{f_{s}\psi_{l}}{Kf_{c}}$, which is a quadratic function about $n$. Further, we have
@@
$\lim_{K\to\infty}\left|[\boldsymbol{\Xi}_{l,n}]_{n,m}\right|\propto\sqrt{K}\delta\left(t_{l,n}-\frac{m}{K}\right).$ (A.4)
@@
From (A.4), when $K\to\infty$, $[\boldsymbol{\Xi}_{\mathrm{D}}]_{n,:}$ can be approximated as a delta function centered at $m=t_{l,n}K$. However, due to the finite sampling size $K$ in practice, we consider the range from $J_{l,e}(n)=(t_{n}-{2}/{K}+1/2)K$ to $J_{l,s}(n)=(t_{n}+{2}/{K}+1/2)K$ to collect the delay indices near the peak effectively. Extending this observation across all $N$ antennas, $\boldsymbol{\Xi}_{\mathrm{D}}$ is characterized as a column-wise block matrix
$\boldsymbol{\Xi}_{l,\mathrm{D}}=\left[\mathbf{0}_{N\times(J_{l,e}-1)},\mathbf{L}_{N\times(J_{l,s}-J_{l,e}+1)},\mathbf{0}_{N\times(K-J_{l,s})}\right],$ (A.5)
where $J_{l,s}=\min_{n}J_{l,s}(n)$ to $J_{l,e}=\max_{n}J_{l,e}(n)$ with $n\in\boldsymbol{\phi}$, and $\mathbf{L}$ is a non-zero matrix.
@@
According to row-wise and column-wise block sparsity of (A.2) and (A.5), we assert that $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ adheres to the following block-sparse structure, with its significant entries localized within the square region $\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\}$, i.e.,
@@
Moreover, based on the definition $\boldsymbol{\Xi}_{l}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$, it follows that, for a fixed column index $k$, $\boldsymbol{\Xi}_{l}$ characterizes the angular distribution corresponding to a specific propagation delay, which is determined by the spatial frequencies across different subcarriers. Accordingly, within the region $\mathcal{A}_{l}$, each column captures the angular response for a given delay, and the width of significant spatial frequency components is approximately determined by the system width and the start spatial frequency, i.e., $i_{l,k}N{f_{s}}/{f_{c}}<I_{l,s}-I_{l,e}$, where $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$ denotes the starting spatial frequency of the $k$-th column. As a result, the nonzero components of $\boldsymbol{\Xi}_{l}$ are confined within a narrow band in each column, leading to a distinct column-wise clustered sparsity pattern.
Appendix B A Proof of Theorem 1
Now, we examine the 2D-IDFT of channel $\mathbf{H}$, i.e., $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$. By leveraging the linearity of the 2D-IDFT, we can focus solely on analyzing the $l$-th path. according to (20), we further define
$\mathbf{H}_{l,p}=\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (B.1)
Applying the shift property of 2D-IDFT, $\mathbf{X}_{l,p}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}_{l,p}\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ should satisfy the block sparsity with its significant entries localized within the square region $\mathcal{B}_{l,p}$ with
$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l,1}+N\epsilon_{l,p},N),\right.$ (B.2)
$k$ $\left.=\mathrm{mod}(k_{l,1}+f_{s}\tau_{l},K),\forall(n_{l,1},k_{l,1})\in\mathcal{A}_{l}\right\}.$
Here, $\mathcal{A}_{l}$ is the non-zero region for $\boldsymbol{\Xi}_{l}$ given by (25), and mod$(a,m)$ is the modulus of $a$ for $m$. Considering all $P$ frequency components, the significant entries of $\mathbf{X}_{l}$ would be localized within the region $\mathcal{B}_{l}\triangleq\cup_{p}\mathcal{B}_{l,p}$. Thus, the proof is completed.